# Deforming Videos to Masks: Flow Matching for Referring Video Segmentation

### v1

@@ Abstract
Referring Video Object Segmentation (RVOS) requires segmenting specific objects in a video guided by a natural language description. The core challenge of RVOS is to anchor abstract linguistic concepts onto a specific set of pixels and continuously segment them through the complex dynamics of a video. Faced with this difficulty, prior work has often decomposed the task into a pragmatic ‘locate-then-segment’ pipeline. However, this cascaded design creates an information bottleneck by simplifying semantics into coarse geometric prompts (e.g, point), and struggles to maintain temporal consistency as the segmenting process is often decoupled from the initial language grounding. To overcome these fundamental limitations, we propose FlowRVS, a novel framework that reconceptualizes RVOS as a conditional continuous flow problem. This allows us to harness the inherent strengths of pretrained T2V models, fine-grained pixel control, text-video semantic alignment, and temporal coherence. Instead of conventional generating from noise to mask or directly predicting mask, we reformulate the task by learning a direct, language-guided deformation from a video’s holistic representation to its target mask. Our one-stage, generative approach achieves new state-of-the-art results across all major RVOS benchmarks. Specifically, achieving a $\mathcal{J}\&\mathcal{F}$ of 51.1 in MeViS (+1.6 over prior SOTA) and 73.3 in the zero shot Ref-DAVIS17 (+2.7), demonstrating the significant potential of modeling video understanding tasks as continuous deformation processes. Code is available at: https://github.com/xmz111/FlowRVS
@@ 1 Introduction
Referring Video Object Segmentation (RVOS) (Khoreva et al., 2018; Gavrilyuk et al., 2018; Hu et al., 2016) requires the machine to segment objects described by natural language queries, which is critical to intelligent systems to precept and interact with the real world (Jiang et al., 2025; Li et al., 2023). The core challenge of RVOS lies in resolving a fundamental spatio-temporal correspondence dilemma: anchoring abstract linguistic concepts onto a dynamic and fine-grained pixel space. Current paradigms tackle this by resorting to approximations: they spatially model correspondence by collapsing semantics into coarse geometric intermediaries like points or bounding boxes, and temporally treat the video as a sequence of independent frames (Lin et al., 2025; Li et al., 2023; Zhu et al., 2024; Bai et al., 2024). This two-stage ‘locate-then-segment’ design creates an inescapable bottleneck with two critical failure points. First, the initial localization stage irreversibly collapses rich linguistic semantics into a coarse geometric prompt like anchors or bounding boxes, forcing the subsequent segmentation model to operate on an impoverished signal stripped of its original reasoning. Second, the ‘temporal segmenting’ stage is often decoupled from the language query, reverting to a generic visual segmenting problem that is inherently fragile to occlusions and unable to leverage the full temporal context of the referring expression (Liang et al., 2025b; Ren et al., 2024).
@@
To address these limitations, we argue that pretrained Text-to-Video (T2V) models offer a fundamental solution since their native capabilities for fine-grained, text-to-pixel synthesis and spatio-temporal reasoning directly counter the bottlenecks of the ‘locate-then-segment’ paradigm. Some previous attempts like VD-IT (Zhu et al., 2024) have successfully repurposed T2V models as powerful feature backbones. In the broader context of dense prediction, DepthFM (Gui et al., 2025) have adapted the entire generative process itself for visual-to-visual tasks like depth estimation based on a image generation model. While these pioneering efforts validate the generative approach, they also expose a shared, critical blind spot: they fail to fully utilize the dynamic, text-driven reasoning that T2V models are capable of and RVOS demands. The feature-extraction approach remains decoupled, forcing a separate decoder to reconstruct temporal relationships from temporally isolated features, squandering the T2V model’s inherent video coherence. Meanwhile, the image-to-depth flows proposed by previous styles completely neglect text condition, rendering them fundamentally incapable of addressing the core RVOS challenge: producing different masks for the same video based on varying textual queries. Thus, we argue that a deeper, more principled alignment with the T2V paradigm is required: one that treats the entire process as a single, unified, language-guided flow from video pixels to required masks.
@@
This unfied flow with existing powerful T2V pretrain model (e.g, Wan) brings several benefits: (1) their pixel-level synthesis training provides fine-grained control, which enables them to distinguish and handle more delicate objects when locating specific targets; (2) their text-condition generation ensures powerful multi-modal alignment, which allows them to ground rich linguistic semantics directly in the pixel space without as much information loss as first mapping in coarse geometric intermediaries; (3) their video-native architecture provides inherent spatio-temporal reasoning, naturally unifying language guidance with temporal consistency.
@@ 2 Related Work
Referring Video Object Segmentation aims to segment a target object within a video based on a natural language expression (Xu et al., 2018; Ding et al., 2023). This task demands both visual-linguistic understanding and robust temporal segmenting. Early approaches often adapted frame-level referring image segmentation models and appended temporal linking mechanisms as a post-processing step (Khoreva et al., 2018). More recent and competitive methods have evolved into more integrated yet predominantly multi-stage pipelines. A dominant paradigm involves a “locate-then-segment” strategy, where a powerful multi-modal model first grounds the textual reference to a spatial region, which then guides a separate segmentation process for each frame.
@@
“Locate-then-Segment” Paradigm manifests in several forms. A significant breakthrough came with the introduction of query-based architectures, inspired by the success of DETR-style (Zhu et al., 2020) transformers in vision tasks (Wu et al., 2022; Yan et al., 2024b), this new paradigm reframed RVOS by treating language as a query to the visual features. Furthermore, multimodal model based approaches like LISA (Lai et al., 2024), VISA (Yan et al., 2024a) and ReferDINO (Liang et al., 2025b; a) leverage a pretrained model’s reasoning or grounding ability, such as LLaVA(Liu et al., 2023) or DETR-based GroundingDINO (Ren et al., 2024), to perform initial object localization, and then introduce a custom-designed mask decoder to generate the final segmentation. A similar philosophy is seen in the “VLM+SAM” family of methods (Luo et al., 2023; Wu et al., 2023), which use a Vision-Language Model for bounding box prediction, followed by a generic segmentation model like SAM (Kirillov et al., 2023; Cuttano et al., 2025) to produce pixel-level masks (He & Ding, 2024; Lin et al., 2025; Pan et al., 2025). Another line of work explores repurposing generative models: VD-IT (Zhu et al., 2024) first extracts features from a pretrained text-to-video diffusion model and then feeds these features into a separate DETR-like architecture for mask prediction. While these methods have pushed the performance boundaries, their multi-stage nature—separating feature extraction from decoding, or localization from segmentation can introduce information bottlenecks and prevent end-to-end optimization of the core correspondence problem.
@@
Generative Modeling is largely catalyzed by the advent of latent diffusion models (Rombach et al., 2022). Building on this success, the frontier rapidly expanded into the temporal domain, leading to a surge of powerful text-to-video (T2V) models (Liu et al., 2024; Wan et al., 2025; Gao et al., 2025). Architectures like Sora have demonstrated a remarkable ability to generate complex, dynamic scenes by scaling up transformer-based models (Peebles & Xie, 2023), proving that these models implicitly learn profound priors about object permanence, motion physics, and the compositional structure of our world. Beyond diffusion models which are predicated on a fixed path to Gaussian noise, Flow Matching (Lipman et al., 2022; Liu et al., 2022) offers a significant theoretical advancement: it learns a velocity field to transport samples along a deterministic ODE path between any two arbitrary distributions. This generalization has been leveraged by pioneering works like DepthFM (Gui et al., 2025), which reformulate tasks like depth estimation as a direct transformation between visual modalities. However, these methods only solve a visual-to-visual translation problem, leaving them unable to address tasks driven by language. Our work makes a critical distinction: we introduce the natural language query as the core conditional force that modulates the entire ODE path. This elevates the framework from a simple translation to a dynamic, multi-modal reasoning engine. It is this specific property that we leverage, proposing to reframe RVOS not as a synthesis task, but as a learned, conditional deformation from the distribution of videos to the distribution of their corresponding masks, instructed by natural language guidance.
@@ 3.3 Analysis of Alternative Paradigms
To motivate our final choices, we first analyze the fundamental limitations of three plausible alternative paradigms for adapting a T2V model to RVOS. As empirically validated in our ablation studies, each of these alternatives fails due to a core mismatch with the nature of RVOS.
@@ 4.1 Benchmark and Metrics
We evaluate our framework on three standard RVOS benchmarks. MeViS (Ding et al., 2023) is a challenging, motion-centric benchmark featuring 2,006 long videos and over 28,000 fine-grained annotations that emphasize complex dynamics. Ref-YouTube-VOS (Wu et al., 2022) is the large-scale benchmark, comprising 3,978 videos that test for generalizability across a wide diversity of objects and scenes. Ref-DAVIS17 (Khoreva et al., 2018) is a high-quality, densely annotated dataset of 90 videos, serving as a key benchmark for segmentation precision and temporal consistency.
@@ 4.2 Implementation Details
Our framework is built upon the publicly available Wan 2.1 text-to-video model, which features a 1.3B parameter Diffusion Transformer (DiT) (Wan et al., 2025). Throughout all training stages, we keep the pretrained text encoder and the VAE encoder entirely frozen. Our training focuses exclusively on fine-tuning the DiT block to learn the conditional flow. Crucially, the VAE decoder is specifically adapted for the segmentation task by being fine-tuned separately on the MeViS training set, allowing it to specialize in reconstructing high-quality masks from the latent space. Our training protocol varies by dataset to align with standard and fair evaluation protocols. For experiments on Ref-YouTube-VOS, we follow a two-stage training strategy. The model is first pre-trained on a combination of static image datasets (RefCOCO/+/g) (Yu et al., 2016; Kazemzadeh et al., 2014) to learn foundational visual-linguistic grounding. Subsequently, this pre-trained model is fine-tuned on the Ref-YouTube-VOS training set. The final weights from this stage are then used for zero-shot evaluation on the Ref-DAVIS17 benchmark without any further fine-tuning to prove FlowRVS’s generalization ability. For the more challenging MeViS dataset, which emphasizes complex motion understanding, we train our model directly on its training set from scratch, without leveraging any static image pre-training. More hyperparameters settings can be found in Appendix A.
@@ 4.4 Ablation Studies
We conduct our ablation studies on the challenging MeViS dataset, as its complex, motion-centric scenarios provide a rigorous testbed for our design choices. To ensure a consistent and fair comparison, all results are reported on the $valid_{u}$ set, following the protocol in prior work (Ding et al., 2023). The results are summarized in Table 2.
@@
(a) MutiStep Noise-to-Mask Flow 32.3 29.6 35.0
(b) Onestep Mask Prediction 36.2 41.5 38.9
@@
Our MutiStep Video-to-Mask Flow
@@
Frozen VAE 29.7 19.6
Add Trainable Conv Head 85.4 53.2
Finetuning Decoder (ours default) 99.1 60.6
@@
Effectiveness of the VAE Adaptation.
A crucial step in our method is adapting the pretrained VAE to accurately transform between the latent space and the pixel space of binary masks. As shown in Table 3, we evaluate several adaptation strategies. In all experiments, we freeze the VAE encoder to preserve its powerful pretrained features and maintain a stable latent space. We then compare three decoder configurations: keeping it frozen, adding a simple convolutional head, and full-parameter finetuning. The results are definitive: fully finetuning the decoder dramatically improves mask reconstruction quality, which directly translates to a significant boost in final RVOS performance. The The visualization results which we provide in Appendix B also demonstrate the effectiveness of such adaptation.
@@ Appendix A Hyperparameters Settings
Appendix B VAE Reconstruction Comparison
As shown in Table 3 and below Figure 4. The vanilla VAE decoder fails to reconstruct accurate masks, a discrepancy we attribute to the extreme domain shift between mask images (binary 0/1) and the natural-photo distribution on which the decoder was pre-trained. Fine-tuning the decoder resolves this failure, indicating that the frozen encoder already encodes sufficient mask-related structure and that only the decoder needs to adapt to the new visual modality.
Figure 4: Visualization of VAE reconstruction results
Appendix C More Qualitative Results

### v2

@@ Abstract
Referring Video Object Segmentation (RVOS) requires segmenting specific objects in a video guided by a natural language description. The core challenge of RVOS is to anchor abstract linguistic concepts onto a specific set of pixels and continuously segment them through the complex dynamics of a video. Faced with this difficulty, prior work has often decomposed the task into a pragmatic ‘locate-then-segment’ pipeline. However, this cascaded design creates an information bottleneck by simplifying semantics into coarse geometric prompts (e.g., point), and struggles to maintain temporal consistency as the segmenting process is often decoupled from the initial language grounding. To overcome these fundamental limitations, we propose FlowRVS, a novel framework that reconceptualizes RVOS as a conditional continuous flow problem. This allows us to harness the inherent strengths of pretrained T2V models, fine-grained pixel control, text-video semantic alignment, and temporal coherence. Instead of conventional generating from noise to mask or directly predicting mask, we reformulate the task by learning a direct, language-guided deformation from a video’s holistic representation to its target mask. Our one-stage, generative approach achieves new state-of-the-art results across all major RVOS benchmarks. Specifically, achieving a $\mathcal{J}\&\mathcal{F}$ of 51.1 in MeViS (+1.6 over prior SOTA) and 73.3 in the zero shot Ref-DAVIS17 (+2.7), demonstrating the significant potential of modeling video understanding tasks as continuous deformation processes. Code is available at: https://github.com/xmz111/FlowRVS
@@ 1 Introduction
Referring Video Object Segmentation (RVOS) (khoreva2018video; gavrilyuk2018actor; hu2016segmentation) requires the machine to segment objects described by natural language queries, which is critical to intelligent systems to precept and interact with the real world (jiang2025affordancesam; li2023locate). The core challenge of RVOS lies in resolving a fundamental spatio-temporal correspondence dilemma: anchoring abstract linguistic concepts onto a dynamic and fine-grained pixel space. Current paradigms often rely on instance-centric approaches, which first identify and then track object instances. While effective, even modern query-based (e.g., ReferFormer (wu2022language)) or VLM-based (e.g., LISA (lai2024lisa)) methods can introduce an information bottleneck by collapsing rich semantics into intermediate object-centric representations. This can limit holistic scene understanding and temporal consistency, especially as the segmentation of each frame, while conditioned, doesn’t stem from a single, unified spatio-temporal deformation process (liang2025referdino; ren2024grounded; bai2024one; lin2025glus).
@@
To address these limitations, we argue that pretrained Text-to-Video (T2V) models offer a fundamental solution since their native capabilities for fine-grained, text-to-pixel synthesis and spatio-temporal reasoning directly counter the bottlenecks of the ‘locate-then-segment’ paradigm. While some attempts use T2V models as powerful frozen feature extractors for a separate decoder (e.g., VD-IT (zhu2024exploring), HCD (zhang2025temporal)), this two-stage design decouples the model’s generative dynamics from the final task. Our work fundamentally differs: we propose to repurpose the entire generative process itself, learning a direct, language-guided deformation flow from video to mask. DepthFM (gui2025depthfm) have adapted the entire generative process itself for visual-to-visual tasks like depth estimation based on a image generation model. While these pioneering efforts validate the generative approach, they also expose a shared, critical blind spot: they fail to fully utilize the dynamic, text-driven reasoning that T2V models are capable of and RVOS demands. The feature-extraction approach remains decoupled, forcing a separate decoder to reconstruct temporal relationships from temporally isolated features, squandering the T2V model’s inherent video coherence. Meanwhile, the image-to-depth flows proposed by previous styles completely neglect text condition, rendering them fundamentally incapable of addressing the core RVOS challenge: producing different masks for the same video based on varying textual queries. Thus, we argue that a deeper, more principled alignment with the T2V paradigm is required: one that treats the entire process as a single, unified, language-guided flow from video pixels to required masks.
@@
This unified flow with existing powerful T2V pretrain model (e.g, Wan) brings several benefits: (1) their pixel-level synthesis training provides fine-grained control, which enables them to distinguish and handle more delicate objects when locating specific targets; (2) their text-condition generation ensures powerful multi-modal alignment, which allows them to ground rich linguistic semantics directly in the pixel space without as much information loss as first mapping in coarse geometric intermediaries; (3) their video-native architecture provides inherent spatio-temporal reasoning, naturally unifying language guidance with temporal consistency.
@@ 2 Related Work
Referring Video Object Segmentation aims to segment a target object within a video based on a natural language expression (xu2018youtube; ding2023mevis). This task demands both visual-linguistic understanding and robust temporal segmenting. Early approaches often adapted frame-level referring image segmentation models and appended temporal linking mechanisms as a post-processing step (khoreva2018video). More recent and competitive methods have evolved into more integrated yet predominantly multi-stage pipelines. A dominant paradigm involves a “locate-then-segment” strategy, where a powerful multi-modal model first grounds the textual reference to a spatial region, which then guides a separate segmentation process for each frame.
@@
“Locate-then-Segment” Paradigm manifests in several forms. A significant breakthrough came with the introduction of query-based architectures, inspired by the success of DETR-style (zhu2020deformable) transformers in vision tasks (wu2022language; yan2024referred), this new paradigm reframed RVOS by treating language as a query to the visual features. Furthermore, multimodal model based approaches like LISA (lai2024lisa), VISA (yan2024visa) and ReferDINO (liang2025referdino; liang2025referdinoplus) leverage a pretrained model’s reasoning or grounding ability, such as LLaVA(liu2023visual) or DETR-based GroundingDINO (ren2024grounded), to perform initial object localization, and then introduce a custom-designed mask decoder to generate the final segmentation. A similar philosophy is seen in the “VLM+SAM” family of methods (luo2023soc; wu2023onlinerefer), which use a Vision-Language Model for bounding box prediction, followed by a generic segmentation model like SAM (kirillov2023segment; cuttano2025samwise) to produce pixel-level masks (he2024decoupling; lin2025glus; pan2025semantic). Another line of work explores repurposing generative models: VD-IT (zhu2024exploring) first extracts features from a pretrained text-to-video diffusion model and then feeds these features into a separate DETR-like architecture for mask prediction. While these methods have pushed the performance boundaries, their reliance on intermediate representations—whether object queries or extracted features—can introduce information bottlenecks that prevent a truly holistic, end-to-end optimization of the video-to-mask correspondence problem.
@@
Generative Modeling is largely catalyzed by the advent of latent diffusion models (rombach2022high). Building on this success, the frontier rapidly expanded into the temporal domain, leading to a surge of powerful text-to-video (T2V) models (liu2024sora; wan2025wan; gao2025seedance). Recent works have begun to leverage these models for RVOS. A notable approach (e.g., VD-IT (zhu2024exploring), HCD (zhang2025temporal)) utilizes T2V models as powerful frozen feature extractors for a separate segmentation decoder. Concurrent work ReferEverything (REM) (bagchi2025refereverythingsegmentingspeakvideos) also explores video-to-mask generation but relies on direct one-step prediction. Our work fundamentally differs: instead of extracting features or learning a static mapping, we repurpose the entire generative process, fine-tuning the core model to learn a direct video-to-mask deformation flow. This avoids the bottleneck inherent in a two-stage pipeline. Furthermore, unlike conditional generation frameworks like ControlNet(zhang2023adding) that add external guidance to a divergent, noise-to-image process, our method learns a convergent, discriminative transformation from the video source itself. Beyond diffusion, Flow Matching (lipman2022flow; liu2022flow) offers a significant theoretical advancement by learning a velocity field to transport samples along a deterministic ODE path. This has been leveraged for visual tasks like depth estimation (e.g., DepthFM (gui2025depthfm)), but these methods are typically text-agnostic. Our work makes a critical distinction: we introduce the natural language query as the core conditional force that modulates the entire ODE path. This elevates the framework from a simple translation to a dynamic, multi-modal reasoning engine, reframing RVOS as a learned, conditional deformation.
@@ 3.3 Analysis of Alternative Paradigms
To motivate our final choices, we first analyze the fundamental limitations of three plausible alternative paradigms for adapting a T2V model to RVOS. As empirically validated in our ablation studies, each of these alternatives fails due to a core mismatch with the nature of RVOS. To ensure a fair comparison, all alternative paradigms were built upon the same Wan2.1 pre-trained model and trained under the exact same supervised setting (same optimizer, learning rate, and duration) as our final model. And we implemented the same fine-tuned VAE to reveal the superiority of our proposed flow-based paradigm.
@@ 4.1 Benchmark and Metrics
We evaluate our framework on three standard RVOS benchmarks. MeViS (ding2023mevis) is a challenging, motion-centric benchmark featuring 2,006 long videos and over 28,000 fine-grained annotations that emphasize complex dynamics. Ref-YouTube-VOS (wu2022language) is the large-scale benchmark, comprising 3,978 videos that test for generalizability across a wide diversity of objects and scenes. Ref-DAVIS17 (khoreva2018video) is a high-quality, densely annotated dataset of 90 videos, serving as a key benchmark for segmentation precision and temporal consistency.
@@ 4.2 Implementation Details
Our framework is built upon the publicly available Wan 2.1 text-to-video model, which features a 1.3B parameter Diffusion Transformer (DiT) (wan2025wan). Throughout all training stages, we keep the pretrained text encoder and the VAE encoder entirely frozen. Our training focuses exclusively on fine-tuning the DiT block to learn the conditional flow. Crucially, the VAE decoder is specifically adapted for the segmentation task by being fine-tuned separately on the MeViS training set, allowing it to specialize in reconstructing high-quality masks from the latent space. Our training protocol varies by dataset to align with the same evaluation protocols of compared methods. For experiments on Ref-YouTube-VOS, we follow a two-stage training strategy. The model is first pre-trained on a combination of static image datasets (RefCOCO/+/g) (yu2016modeling; kazemzadeh2014referitgame) to learn foundational visual-linguistic grounding. Subsequently, this pre-trained model is fine-tuned on the Ref-YouTube-VOS training set. The final weights from this stage are then used for zero-shot evaluation on the Ref-DAVIS17 benchmark without any further fine-tuning to prove FlowRVS’s generalization ability. For the more challenging MeViS dataset, which emphasizes complex motion understanding, we train our model directly on its training set from scratch, without leveraging any static image pre-training. More hyperparameters settings can be found in Appendix A.
@@ 4.4 Ablation Studies
We conduct our ablation studies on the challenging MeViS dataset, as its complex, motion-centric scenarios provide a rigorous testbed for our design choices. To ensure a consistent and fair comparison, all results are reported on the $valid_{u}$ set, following the protocol in prior work (ding2023mevis). The results are summarized in Table 2.
@@
(a) MultiStep Noise-to-Mask Flow 32.3 29.6 35.0
(b) Onestep Mask Prediction 38.9 36.2 41.5
@@
Our MultiStep Video-to-Mask Flow
@@
Frozen VAE 98.8 60.0
Finetuning Decoder (ours default) 99.1 60.9
@@
VAE Adaptation and frozen VAE.
A crucial step in our method is adapting the pretrained VAE to accurately transform between the latent space and the pixel space of binary masks. As shown in Table 3, we evaluate the effectiveness of the decoder adaptation. Consistent with observations in concurrent work REM (bagchi2025refereverythingsegmentingspeakvideos), we find that the frozen VAE decoder can already support competitive segmentation performance (60.0 $\mathcal{J}\&\mathcal{F}$). However, fully finetuning the decoder further bridges the domain gap between continuous video latents and binary masks. This adaptation yields superior reconstruction quality and provides a solid performance gain (+0.9 $\mathcal{J}\&\mathcal{F}$) over the frozen baseline.
@@ 5 Conclusion and Future Work
6 Acknowledgment
This work was supported by Zhejiang Province Natural Science Foundation of China under Grant No. LQN25F030008 and the National Natural Science Foundation of China under Grant No. 62403429.
Ethics Statement
This manuscript does not cover any topics or experiments related to ethics.
Reproducibility Statement
This manuscript provides detailed implement details, hyperparameter settings for reproduction. In addition, the code will be open-sourced later at depended on acceptance
Use of LLMs
We leverage large language models (LLMs) to draft and polish our manuscripts, only for ensuring each sentence is clearer and the narrative flows coherently.
@@ Appendix A Hyperparameters Settings
Appendix B More Qualitative Results
@@
Appendix C Comprehensive Results
Figure 7: failure cases of FlowRVS results.
Figure 8: novel action phrases out of MeViS training.
While our method demonstrates strong performance, it is not without limitations. We present two typical failure modes on Figure 7. For complex relational phrases requiring fine-grained interaction understanding, such as ”swings its tail and strikes the head” on the top row, the model correctly identifies the primary subject (the horse) but fails to isolate the specific horse performing the action. This suggests a limitation in comprehending intricate multi-part actions. In scenarios with multiple similar objects and temporal ordering cues (”the second bird”) on the bottom row, the model struggles to accurately resolve the ambiguity. It incorrectly segments the first bird that moves, indicating that while our temporal modeling is strong, it can be confounded by challenging counting and ordering logic within dense scenes.
To address the question of generalization to unseen actions and challenging description, we tested our model on phrases not present in the training data, such as ”the dog somersaulting.” As shown on Figure 8, FlowRVS successfully identifies and segments the dog throughout its complex, non-rigid motion. This demonstrates that our method does not merely memorize action-object pairings from the training set. Instead, by leveraging the rich spatio-temporal and semantic priors from the pretrained T2V model, it develops a more fundamental understanding based on open-set vocabulary queries that allows it to generalize to novel and dynamic actions.
Appendix D Details about proposed improvements
D.1 Boundary-Biased Sampling (BBS)
BBS is a training strategy to emphasize the crucial initial step of the flow ($t=0$). Formally, we sample the timestep $t$ from a mixed probability distribution, whose probability density function $f(t)$ is defined as: $f(t)=p\cdot\delta(t)+(1-p)\cdot\mathcal{U}(t|0,1)$ where $\delta(t)$ is the Dirac delta function representing a point mass at $t=0$, $\mathcal{U}(t|0,1)$ is the uniform distribution on the interval $[0,1]$, and $p$ is the bias probability. In practice, this means we sample $t=0$ with probability (p=0.5).
D.2 Start-Point Augmentation (SPA)
SPA is a crucial regularizer to improve the robustness of our convergent (video-to-mask) flow. For a given video $V$, the VAE encoder $E_{\phi}$ predicts a posterior distribution $q(z|V)=\mathcal{N}(z|\bm{\mu}_{V},\bm{\sigma}^{2}_{V})$, where $\bm{\mu}_{V}$ and $\bm{\sigma}^{2}_{V}$ are directly inherited from the VAE. Instead of using the deterministic mean $\bm{\mu}_{V}$, SPA samples the starting point $\bm{z^{\prime}}_{0}$ from this posterior: $\bm{z^{\prime}}_{0}\sim\mathcal{N}(\bm{z}|\bm{\mu}_{V},\bm{\sigma}^{2}_{V})$. This sampled latent $\bm{z^{\prime}}_{0}$ is then normalized before being used the ODE solver. By augmenting the training data with samples from the local neighborhood of each video’s true latent representation, SPA forces the model to learn a smoother and more generalizable velocity field.
D.3 Direct Video Injection (DVI)
DVI provides the model with a persistent anchor to the original video content throughout the ODE trajectory. It is implemented by concatenating the current state $\bm{z}_{t}$ with the initial video latent $\bm{z}_{0}$. This changes the input tensor’s shape from $[B,C,T,H,W]$ to [B,2C,T,H,W][B,2*C,T,H,W]. This is handled by modifying the first convolutional layer of the DiT to accept 2C2*C input channels, while all subsequent layers remain unchanged.